HID- 105891775offset-prev <Chapter 19-Homework Problem 19.52 Cakculate the resistance of R Constants Express your answer in ohms to three significant figures Referance the circuit shown below. (Assume Vi-100V and V 170V) 40Ω Re i 10 A 20 n Submit Part B Calcuilate the corent hrough the 40 esistor ssue that the batteries have no intermal Express your answe i amperes to three significant figures. Y Part G
Answer #1

here,

V1 = 100 V

V2 = 170 V

a)

for the left small loop,

V1 = i * R

100 = 10 * R

R = 10 ohm

b)

let the current in 40 ohm resistance be i3

for the middle loop,

V2 = 40 * i3

170 = 40*i3

i3 = 4.25 A

the current through 40 ohm resistance is 4.25 A